# Model-based Manipulation of Deformable Objects with Non-negligible Dynamics as Shape Regulation

### v1

@@ Abstract
Model-based manipulation of deformable objects has traditionally dealt with objects in the quasi-static regimes, either because they are extremely lightweight/small or constrained to move very slowly. On the contrary, soft robotic research has made considerable strides toward general modeling and control - despite soft robots and deformable linear objects being very similar from a mechanical standpoint. In this work, we leverage these recent results to develop a fully dynamic framework of slender deformable objects grasped at one of their ends by a robotic manipulator. We introduce a dynamic model of this system using functional strain parameterizations and describe the manipulation challenge as a regulation control problem. This enables us to define a fully model-based control architecture, for which we can prove analytically closed-loop stability and provide sufficient conditions for steady state convergence to the desired manipulation state. The nature of this work is intended to be markedly experimental. We propose an extensive experimental validation of the proposed ideas. For that, we use a 7-DoF robot tasked with the goal of positioning the distal end of six different electric cables, moving on a plane, in a given position and orientation in space.
Note: a video showcasing the performance of the proposed modeling and control strategies is available here.
@@ I Introduction
This is the case also for the simplest object geometries that we will be the focus of the present work: slender or deformable linear [^1] It is worth stressing that linear does not refer to the dynamics or kinematics of these systems being linear, but to the fact that these objects look like a curved line. In fact, their dynamics is highly nonlinear, as we will discuss later in the paper. To avoid confusion, we sometimes employ the less common denomination of slender object. objects (DLOs). These are deformable bodies whose spatial configuration can be largely specified in one dominant spatial dimension. A substantial body of literature focuses on very lightweight, small, or slowly-moving DLOs (i.e., quasi-static regimes). These assumptions allow circumventing the challenge of formulating their dynamics by focusing on purely kinematic descriptions [9, 10, 11, 12, 13], usually relying on the deformation Jacobian paradigm and shape servoing. Machine learning-based variations have also been considered [14, 12]. Learning strategies are nowadays a common choice when it comes to relaxing the quasi-static hypothesis to move into the dynamic manipulation realm [15, 16, 17, 18, 19, 20, 21, 22]. However, learning-based control comes with its own well-known limitations. To name a few: lack of guarantees, need for large amounts of data, and potential unsafety of the learning process.
@@
Figure 1: Pictorial representation of the task that we investigate in this work: the manipulation of slender deformable objects via a generic manipulator holding them to one of their ends by its end effector. We superimpose the reference frames important to describe the dynamics of the slender object.
@@
This work aims to introduce and experimentally validate a fully model-based perspective on manipulating slender deformable objects that does not require applying any quasi-static assumption on the object’s behavior. To achieve this goal, we put ourselves in the context of classical control of deformable mechanical systems [23, 24, 25, 26, 27]. More precisely, we borrow from recent advancements in model-based control of soft robots [28], which are systems entirely made of continuously deformable soft materials [29, 30]. The past few years have seen an exponential increase in control-oriented models of soft robots [31, 32]. These include shape [^2] Their kinematic version have been independently investigated in deformable objects literature [33, 34, 35]. [36, 37] and - more recently - strain parametrizations [38, 39, 40, 41]. In turn, the availability of finite dimensional interpretable models has driven the flourishing of model-based control techniques [42, 43, 44, 45, 46, 47, 48, 49, 50, 51] that combine improved performance with theoretical guarantees.
@@
To conclude, we propose a dynamic modeling framework for DLOs that have been pre-grasped at one of their ends by a generic robotic manipulator (see Fig. 1). The framework uses strain-based functional parameterization to model DLOs with a non-negligible dynamic response. We then cast the manipulation challenge as a regulation control problem and introduce a control architecture (see Fig. 2), for which we analytically prove closed-loop stability.
@@ II Strain-Based Dynamic Modeling of DLOs
Our goal is to introduce a fully dynamic model of a robot-object pair as described by Fig. 1. We start by introducing a dynamic model of the object - as it is the part that requires most of the attention - and only at the end we combine it with the model of the robot.
@@ II-A Additional assumptions
Here, we consider a few assumptions for the sake of simplicity of derivations but without loss of generality. We considered the object inextensible and not experiencing shear strain. We focus on the polynomial curvature functional parameterization [53], which is representative of the general strain case discussed in [39]. Moreover, we propose derivations in the case of planar motions of the object, as this is the case we consider in our experimental validation, and - again - including full motions in 3D only makes derivations more complex without any substantial shift in the core ideas introduced here [54, 55]. Furthermore, the control algorithms that we propose later in this manuscript are agnostic to these simplifying assumptions and can be applied to the general case without modifications.
@@ II-B Object’s Backbone Kinematics
where $L$ is the total length of the object, and $x,y\in\mathbb{R}$ are two extra parameters. It is immediate to verify that $(x_{\mathrm{B}},y_{\mathrm{B}},\alpha_{\mathrm{B}})(0)=(x,y,\phi)$ thus giving an immediate interpretation to the three parameters as the configuration of the robot’s end effector.
@@
Figure 2: Block diagram outline of the control strategy
@@ II-D Object’s Dynamics
The direction of the gravitational field is considered directed along $-y$ in the $\{S_{\mathrm{B}}\}$ frame.
@@
$\small\begin{split}B_{\mathrm{O}}(q_{\mathrm{O}})\ddot{q}_{\mathrm{O}}&+C_{\mathrm{O}}(q_{\mathrm{O}},\dot{q}_{\mathrm{O}})\dot{q}_{\mathrm{O}}+G_{\mathrm{O}}(q_{\mathrm{O}})\\ +&\begin{bmatrix}kH&0_{n+1\times 3}\\ 0_{3\times n+1}&0_{3\times 3}\end{bmatrix}\begin{bmatrix}\Theta-\bar{\Theta}\\ x\\ y\\ \phi\end{bmatrix}\\ +&\begin{bmatrix}\beta H&0_{n+1\times 3}\\ 0_{3\times n+1}&0_{3\times 3}\end{bmatrix}\dot{q}_{\mathrm{O}}=\begin{bmatrix}0_{n\times 1}\\ F_{\mathrm{x}}\\ F_{\mathrm{y}}\\ \tau_{\phi}\end{bmatrix},\end{split}$ (8)
where $F_{\mathrm{x}}$, $F_{\mathrm{y}}$, and $\tau_{\phi}$ are generalized forces (two forces and torque) representing the action of the robot on the object, $H$ is the $n\times n$ Hankel matrix, and the internal stiffness and damping coefficients of the object averaged along its length are $k$ and $\beta$.
@@ II-E Complete Dynamical Model
$\small\begin{split}&\begin{bmatrix}B_{\mathrm{r},\mathrm{r}}(q_{\mathrm{r}},\Theta)&B_{\mathrm{r},\Theta}(q_{\mathrm{r}},\Theta)\\ B_{\mathrm{r},\Theta}^{T}(q_{\mathrm{r}},\Theta)&B_{\Theta,\Theta}(\Theta)\end{bmatrix}\begin{bmatrix}\ddot{q}_{\mathrm{r}}\\ \ddot{\Theta}\end{bmatrix}+C(q_{\mathrm{r}},\Theta,\dot{q}_{\mathrm{r}},\dot{\Theta})\begin{bmatrix}\dot{q}_{\mathrm{r}}\\ \dot{\Theta}\end{bmatrix}\\ &+\begin{bmatrix}0_{n_{\mathrm{r}}\times 3}&0_{n+1\times n_{\mathrm{r}}}\\ 0_{n_{\mathrm{r}}\times n+1}&\beta H\end{bmatrix}\begin{bmatrix}\dot{q}_{\mathrm{r}}\\ \dot{\Theta}\end{bmatrix}+\begin{bmatrix}G_{\mathrm{r}}(q_{\mathrm{r}},\Theta)\\ G_{\Theta}(q_{\mathrm{r}},\Theta)\end{bmatrix}\\ &+\begin{bmatrix}0_{n_{\mathrm{r}}\times 3}&0_{n+1\times n_{\mathrm{r}}}\\ 0_{n_{\mathrm{r}}\times n+1}&kH\end{bmatrix}\begin{bmatrix}{q}_{\mathrm{r}}\\ (\Theta-\bar{\Theta})\end{bmatrix}=\begin{bmatrix}\tau\\ 0_{n+1\times 1}\end{bmatrix}.\end{split}$ (9)
@@ III-B Low-level Control
Let’s consider the case in which the control inputs $\tau$ are used to regulate the collocated configuration variables $q_{\mathrm{r}}$.
@@ III-B1 Stability of the zero dynamics
Where $\phi^{*}$ is the end effector orientation associated with the fixed robot configuration $q^{*}_{\mathrm{r}}$. This collocated zero dynamics is such that
@@
if $\beta>0$. Moreover, the ${\Theta}^{*}$ such that
$G_{\Theta}(\Theta^{*},\phi^{*})+kH{(\Theta^{*}-\bar{\Theta})}=0,$ (14)
is an asymptotically stable equilibrium of the zero dynamics if the potential energy is convex in $\Theta^{*}$ (i.e., $k$ large enough). The stability is global if the potential energy is globally convex.
To prove these statements it is sufficient to take the object’s total energy as a Lyapunov candidate and follow similar steps as in the soft robotic survey paper [28], that we do not report here for the sake of space.
@@ III-B2 Control strategy
$\tau=K_{\mathrm{P}}(q^{*}_{\mathrm{r}}-q_{\mathrm{r}})+K_{\mathrm{D}}(-\dot{q}_{\mathrm{r}})+G_{\mathrm{r}}(q^{*}_{\mathrm{r}},\Theta),$ (15)
where $K_{\mathrm{P}},K_{\mathrm{D}}$ are two positive definite control gain matrices. The only condition necessary for the closed loop stability is that $\beta>0$. No condition on $k$ is required. Again, for the sake of conciseness, we do not report the stability proof that can be obtained by recasting the steps of the one we propose in the soft robotic control paper [59].
Remarkably, this simple control strategy translates into the standard PD with gravity compensation when imposing the following hypothesis
$\frac{\partial G_{\mathrm{r}}}{\partial\Theta}\simeq 0,$ (16)
which is often verified in practice as it reflects the common case of the mass of the object being considerably smaller than the robot’s mass.
@@ IV-A1 Test Objects
Figure 3: The test objects OB1-OB6 from left to right.
The objects considered here are assumed to have constant density and cross-section along their length; however, instead of using the exact continuous distribution of $\rho(s,d)$ when evaluating (6) and (7), we simplify using an approximation of six discrete lump masses. These are equally spaced along the length of the object at positions $s=\{\frac{1}{12},\frac{3}{12},\frac{5}{12},\frac{7}{12},\frac{9}{12},\frac{11}{12}\}$, each with a mass of $\frac{m_{L}}{6}$ distributed across the total diameter. Different masses $m_{0}$ and $m_{1}$ are also included at $s=0$ and $s=1$, to account for separate masses concentrated at the base and end of the object. In terms of unmodeled non-idealities, approximately 0.05-0.1m at the end of the objects is rigid to facilitate the attachment of the endpoint mass. Table I summarises the properties of the objects.

### v2

# Controlling Deformable Objects with Non-negligible Dynamics: a Shape-Regulation Approach to End-Point Positioning

@@ Abstract
Model-based manipulation of deformable objects has traditionally dealt with objects while neglecting their dynamics, thus mostly focusing on very lightweight objects at steady state. At the same time, soft robotic research has made considerable strides toward general modeling and control, despite soft robots and deformable objects being very similar from a mechanical standpoint. In this work, we leverage these recent results to develop a control-oriented, fully dynamic framework of slender deformable objects grasped at one end by a robotic manipulator. We introduce a dynamic model of this system using functional strain parameterizations and describe the manipulation challenge as a regulation control problem. This enables us to define a fully model-based control architecture, for which we can prove analytically closed-loop stability and provide sufficient conditions for steady state convergence to the desired state. The nature of this work is intended to be markedly experimental. We provide an extensive experimental validation of the proposed ideas, tasking a robot arm with controlling the distal end of six different cables, in a given planar position and orientation in space.
@@ I Introduction
This is the case also for the simplest object geometries that we will be the focus of the present work: slender or deformable linear [^1] It is worth stressing that linear does not refer to the dynamics or kinematics of these systems being linear, but to the fact that these objects look like a curved line. In fact, their dynamics are highly nonlinear, as we will discuss later in the paper. To avoid confusion, we sometimes employ the less common denomination of slender object. objects (DLOs). These are deformable bodies whose spatial configuration can be largely specified in one dominant spatial dimension.
@@
Figure 1: Pictorial representation of the task that we investigate in this work: the manipulation of slender deformable objects via a generic manipulator holding them at one of their ends by its end effector. We superimpose the important reference frames.
A substantial body of literature focuses on very lightweight, small, or slowly-moving DLOs (i.e., quasi-static regimes). These assumptions allow circumventing the challenge of formulating their dynamics by focusing on purely kinematic descriptions [9, 10, 11, 12, 13], usually relying on the deformation Jacobian paradigm and shape servoing. Interestingly, [14] extends differential geometry to model linear object deformation—including flexure, torsion, and extension—for grasping and path planning. However, it remains limited to static cases, where inertia can be neglected. Machine learning-based variations of this general approach have also been considered [15, 12]. Learning strategies are nowadays a common choice when it comes to relaxing the quasi-static hypothesis to move into the dynamic manipulation realm [16, 17, 18, 19, 20, 21, 22]. However, learning-based control comes with its own well-known limitations. Some of them are the lack of guarantees, the need for large amounts of data, and the potential safety concerns of the learning process.
@@
This work aims to introduce and experimentally validate a fully control-oriented model-based perspective on manipulating slender deformable objects that does not require applying any quasi-static assumption on the object’s behavior. To achieve this goal, we put ourselves in the context of classical control of deformable mechanical systems [14, 23, 24, 25, 26, 27, 28]. More precisely, we borrow from recent advancements in model-based control of soft robots [29], which are systems entirely made of continuously deformable soft materials [30, 31]. The past few years have seen an exponential increase in control-oriented models of soft robots [32, 33]. These include shape [^2] Their kinematic version have been independently investigated in deformable objects literature [34, 35, 36]. [37, 38] and - more recently - strain parametrizations [39, 40, 41, 42]. In turn, the availability of finite-dimensional interpretable models has driven the flourishing of model-based control techniques [43, 44, 45, 46, 47, 48, 49, 50, 51, 52] that combine improved performance with theoretical guarantees.
@@
To conclude, we propose a dynamic modeling framework for slender objects that have been pre-grasped at one of their ends by a generic robotic manipulator (see Fig. 1). The framework uses strain-based functional parameterization to model DLOs with a non-negligible dynamic response. We then cast the manipulation challenge as a regulation control problem and introduce a control architecture (see Fig. 2), for which we analytically prove closed-loop stability. Compared to recent works [53, 54, 55, 56] proposing control schemes within a dynamic context for single- and dual-arm DLO manipulation, our work: 1) proves closed-loop stability with explicit conditions, 2) models coupled robot-object dynamics with stability guarantees, and 3) includes gravity effects, analyzing behavior in gravity-influenced planes.
@@ II Strain-Based Dynamic Modeling of DLOs
Our goal is to introduce a fully dynamic model of a robot-object pair as described by Fig. 1. We start by introducing a dynamic model of the object, as it is the part that requires attention, and only at the end do we combine it with the model of the robot.
@@ II-A Additional assumptions
Here, we consider a few assumptions for the sake of simplicity of derivations but without loss of generality. We considered the object inextensible and not experiencing shear strain in the plane. We also neglect to model any out-of-plane offset curvature and torsional strain, as the gravitational forces dominate in the range of end-effector orientations we validate the model in. We focus on the polynomial curvature functional parameterization [58], which is representative of the general strain case discussed in [40]. Moreover, we propose derivations in the case of planar motions of the object, as this is the case we consider in our experimental validation, and - again - including full motions in 3D only makes derivations more complex without any substantial shift in the core ideas introduced here [59, 60]. Furthermore, the control algorithms that we propose later in this manuscript are agnostic to these simplifying assumptions and can be applied to the general case without modifications.
@@ II-B Object’s Backbone Kinematics
where $L$ is the total length of the object, and $x,y\in\mathbb{R}$ are two extra parameters. It is immediate to verify that $(x_{\mathrm{B}},y_{\mathrm{B}},\alpha_{\mathrm{B}})(0)=(x,y,\pi+\phi)$ thus giving an immediate interpretation to the three parameters as the configuration of the robot’s end effector.
@@ II-D Object’s Dynamics
The direction of the gravitational field is considered directed along y-y in the $\{S_{\mathrm{B}}\}$ frame.
@@
[…]
where $F_{\mathrm{x}}$, $F_{\mathrm{y}}$, and $\tau_{\phi}$ are generalized forces (two forces and torque) representing the action of the robot on the object, $H$ is the $n\times n$ Hankel matrix, and the internal stiffness and damping coefficients of the object averaged along its length are $k$ and $\beta\succ 0$.
@@ II-E Complete Dynamical Model
[…]
@@
Figure 2: A block diagram outline of the control strategy. The input is a task goal $\zeta^{*}$ which can be achieved by regulation of the object’s shape. We use a nonlinear model-constrained optimization to map this to a Cartesian object base pose $(x^{*},y^{*},\phi^{*})$, providing a control input to the coupled manipulator-object system.
@@ III-B Low-level Control
At this stage, we use the control inputs $\tau$ to regulate the collocated configuration variables $q_{\mathrm{r}}$. Our goal for this section is to analyze the closed-loop control problem and propose a simple but provably stable control strategy. This layer will then be used by the high-level controller to ultimately solve the control goal stated above.
@@ III-B1 Stability of the zero dynamics
Where $\phi^{*}$ is the end effector orientation associated with the fixed robot configuration $q^{*}_{\mathrm{r}}$. This equation describes the behavior of the object when the robot is fixed in some configuration, and analyzing its stability is a fundamental step in designing a robot-side controller and assessing its performance. We do this in the following proposition.
Proposition 1
The state $(\Theta^{*},0)\in\mathbb{R}^{2n}$ is an asymptotically stable equilibrium of (12) if an open neighborhood $\mathcal{N}(\Theta)$ $\subseteq\mathbb{R}^{2}$ of $\Theta^{*}$ exists such that $\forall\Theta\in\mathcal{N}(\Theta^{*})/\{\Theta^{*}\}$
$\begin{split}(U_{G_{\Theta}}(\Theta,\phi^{*})+&U_{K}(\Theta))>(U_{G_{\Theta}}(\Theta^{*},\phi^{*})+U_{K}(\Theta^{*}))\\ +&\left(\frac{\partial}{\partial\Theta}(U_{G_{\Theta}}(\Theta,\phi^{*})+U_{K}(\Theta))\right)\bigg{|}^{T}_{\Theta=\Theta^{*}}(\Theta-\Theta^{*})\end{split}$ (13)
where $U_{G_{\Theta}}$ and $U_{K}$ represent the potential energy associated with gravity and elasticity of the DLO, respectively.
Proof 1
Taking inspiration from the soft robotics literature, we propose to use a modified version of the Lyapunov function in [29, Theorem 1], which is
$\begin{split}V=&\frac{1}{2}\dot{\Theta}^{T}B_{\Theta,\Theta}\dot{\Theta}+U_{G_{\Theta}}(\Theta,\phi^{*})-U_{G_{\Theta}}(\Theta^{*},\phi^{*})+U_{K}(\Theta)\\ -&U_{K}(\Theta^{*})+\left(kH{(\Theta^{*}-\bar{\Theta})}+G_{\Theta}(\Theta^{*},\phi^{*})\right)(\Theta^{*}-\Theta).\end{split}$ (14)
Taking into account condition (13) and the fact that the kinetic energy is always strictly positive definite in $\dot{\Theta}$, it follows that $V$ is positive definite. Next, we analyze the sign of its time derivative, which is
$\begin{split}\dot{V}=&\frac{1}{2}\dot{\Theta}^{T}\dot{B}_{\Theta,\Theta}\dot{\Theta}+\dot{\Theta}^{T}B_{\Theta,\Theta}\ddot{\Theta}+\dot{\Theta}^{T}(kH{(\Theta-\bar{\Theta})}+G_{\Theta}(\Theta,\phi^{*}))\\ -&\dot{\Theta}^{T}(kH{(\Theta^{*}-\bar{\Theta})}+G_{\Theta}(\Theta^{*},\phi^{*}))\\ =&\frac{1}{2}\dot{\Theta}^{T}\dot{B}_{\Theta,\Theta}\dot{\Theta}+\dot{\Theta}^{T}(kH{(\Theta-\bar{\Theta})}+G_{\Theta}(\Theta,\phi^{*}))\\ &+\dot{\Theta}^{T}\left(-C_{\Theta,\Theta}\dot{\Theta}-kH{(\Theta-\bar{\Theta})}-G_{\Theta}(\Theta,\phi^{*})-\beta H\dot{\Theta}\right)\\ =&-\dot{\Theta}^{T}\beta H\dot{\Theta}\leq 0,\end{split}$ (15)
where we leverage the system’s passivity, as expressed by $\dot{B}_{\Theta,\Theta}-2C_{\Theta,\Theta}=0$, and the equilibrium condition $kH{(\Theta^{*}-\bar{\Theta})}+G_{\Theta}(\Theta^{*},\phi^{*})=0$. Consequently, equation (15) is shown to be negative semidefinite, but thanks to LaSalle’s principle, we can guarantee that the system converges to the set $(\Theta^{*},0)$, provided that the equilibrium configuration $\Theta^{*}$ is the only configuration in $\mathcal{N}(\Theta^{*})$ satisfying $\ddot{\Theta}=0$ for $\dot{\Theta}=0$. $\square$
Then, the collocated zero dynamics (12) is such that
@@
if $\beta>0$.
@@ III-B2 Control strategy
$\tau=K_{\mathrm{P}}(q^{*}_{\mathrm{r}}-q_{\mathrm{r}})+K_{\mathrm{D}}(-\dot{q}_{\mathrm{r}})+G_{\mathrm{r}}(q_{\mathrm{r}},\Theta).$ (17)
where $K_{\mathrm{P}},K_{\mathrm{D}}\succ 0$ are two positive definite control gain matrices. The only condition necessary for the closed loop stability is that $\beta>0$.
Proposition 2
There always exist a large-enough proportional gain $K_{\mathrm{P}}$ such that the trajectories of the closed-loop system (17), (9) are bounded and converge asymptotically to the equilibrium state $(q_{r}^{*},\Theta^{*},0,0)$, where $\Theta^{*}$ is a solution of
$G_{\Theta}(\Theta^{*},\phi^{*})+kH{(\Theta^{*}-\bar{\Theta})}=0.$ (18)
Proof 2
The proof follows directly from Proposition 1, following the same steps of the proof of Theorem 1 in the soft robotics control paper [63]. $\square$
Furthermore, we can provide even stronger convergence properties for the closed loop under slightly stronger hypotheses, as discussed in the following proposition.
Proposition 3
The closed loop (17), (9) converges asymptotically to the unique solution of (18) in the region of attraction $\mathcal{N}$, if
$\left(kH+\frac{\partial G_{\Theta}(\Theta,\phi^{*})}{\partial\Theta}\right)\succ 0,\quad\forall\Theta\in\mathcal{N}.$ (19)
Proof 3
The statement has two parts. The first one is that the solution in $\mathcal{N}$ is unique. This follows by the application of Hadamard’s global inverse function theorem since the left-hand side of (19) is the Jacobian of the left-hand side of (18). Also, the left-hand side of (19) is proper because it is radially unbounded as $\|\Theta\|\to\infty$, the linear term $kH(\Theta-\bar{\Theta})$ dominates because it grows without bound. This allows to state the property globally if $\mathcal{N}$ coincides with the whole configuration space.
The second part is about asymptotic convergence. This one can be proven by relying on a variation of the Lyapunov candidate that we proposed in the proof of Proposition 1
$\begin{split}V&=\frac{1}{2}\begin{bmatrix}\dot{q}_{\mathrm{r}}\\ \dot{\Theta}\end{bmatrix}^{\top}\begin{bmatrix}B_{\mathrm{r},\mathrm{r}}(q_{\mathrm{r}},\Theta)&B_{\mathrm{r},\Theta}(q_{\mathrm{r}},\Theta)\\ B_{\mathrm{r},\Theta}^{T}(q_{\mathrm{r}},\Theta)&B_{\Theta,\Theta}(\Theta)\end{bmatrix}\begin{bmatrix}\dot{q}_{\mathrm{r}}\\ \dot{\Theta}\end{bmatrix}\\ &+U_{G_{\Theta}}(\Theta,\phi)-U_{G_{\Theta}}(\Theta^{*},\phi^{*})+U_{K}(\Theta)-U_{K}(\Theta^{*})\\ &+\left(kH{(\Theta^{*}-\bar{\Theta})}+G_{\Theta}(\Theta^{*},\phi^{*})\right)(\Theta^{*}-\Theta)\\ &+\frac{1}{2}(q^{*}_{r}-q_{r})^{T}K_{p}(q^{*}_{r}-q_{r}).\end{split}$ (20)
The proof then follows similar steps, with the only addition of noticing that (19) implies (13). $\square$
In the soft robotics literature, (19) is called elastic dominance, and its goal is to ensure that the physical compliance of the robot (or, in this case, of the object) is high-enough when compared to gravity.
Remark 1
[…]
@@ IV-A1 Test Objects
Figure 3: The test objects used in the experiments, OB1-OB6 from left to right. These are high-voltage electric cables used in the context of the electric car industry. The cables end with a plug connector.
The objects considered here are assumed to have constant density and cross-section along their length; however, instead of using the exact continuous distribution of $\rho(s,d)$ when evaluating (6) and (7), we simplify using an approximation of six discrete lump masses. These are equally spaced along the length of the object at positions $s=\{\frac{1}{12},\frac{3}{12},\frac{5}{12},\frac{7}{12},\frac{9}{12},\frac{11}{12}\}$, each with a mass of $\frac{m_{\mathrm{L}}}{6}$ distributed across the total diameter. Different masses $m_{0}$ and $m_{1}$ are also included at $s=0$ and $s=1$, to account for separate masses concentrated at the base and end of the object. In terms of unmodeled non-idealities, approximately 0.05-0.1m at the end of the objects is rigid to facilitate the attachment of the endpoint mass. Table I summarises the properties of the objects.